# Interpretable User Satisfaction Estimation for Conversational Systems with Large Language Models

### v1

@@ 4 Evaluation
Acc
P
R
F1
Acc
P
R
F1
Acc
P
R
F1
Acc
P
R
F1
Lin-ada
66.5
67.1
66.5
66.8
27.6
44.6
27.6
31.2
38.2
53.2
38.2
42.9
28.4
48.6
28.4
33.9
USDA
69.5
34.8
50.0
41.0
47.8
22.9
47.8
30.9
70.6
68.3
70.6
64.9
38.6
70.0
38.6
24.0
ASAP
69.5
48.3
69.5
57.0
57.1
57.9
57.1
56.9
65.8
67.3
65.8
66.4
70.2
69.5
70.2
67.6
Reward
34.6
11.9
34.6
17.8
48.5
23.5
48.5
31.7
61.9
38.3
61.9
47.3
58.1
33.7
58.1
42.6
ZeroShot
76.6
75.7
76.6
74.1
57.3
70.4
57.3
61.7
75.2
76.9
75.2
75.6
63.2
40.0
63.2
49.0
FewShot
69.4
57.9
69.4
57.2
50.0
65.6
50.0
44.2
68.2
46.8
68.2
55.5
62.2
40.0
62.2
48.7
SPUR 78.4 77.5 78.4 77.4 64.4 66.1 64.4 63.2 76.6 76.8 76.6
75.2
70.5 70.5 70.5 70.0
USE under Few-Shot Setting. Table 1 shows the performance of each model trained with a small number of training examples. The performance metrics are weighted based on the label distributions due to the data imbalance in the different datasets. The training set sizes are shown beside the name of each dataset and the remaining $80\%$ of the data is used for testing. The number of items in the satisfaction and dissatisfaction rubrics is ten, respectively. Three task-oriented datasets have larger training sizes because we want to ensure that there are at least ten conversations with satisfaction label and ten conversations with dissatisfaction label to derive SPUR’s rubrics.
Comparing the performance between ZeroShot and SPUR in Table 1, the effectiveness of the rubrics can be observed. Prompting with learned rubrics can provide better guidance for LLMs than prompting with a handcrafted set of features Bodigutla et al. (2020). On the other hand, FewShot has worse performance compared to other methods because the examples provided in the prompt cannot cover many types of satisfaction/dissatisfaction conversational patterns, and the decision is usually biased by the examples provided in the prompt.
The performance of the Reward model (reward deberta, ) validates our hypothesis that Reward models used for RLHF cannot be a proxy for USE. Because Reward models are usually trained with auxiliary human feedback, this reward is not learned from the perspective of the user who was involved in the conversation with the AI agent (Kirk et al., 2023).
Embedding methods perform worse than SPUR in Table 1. This is not unexpected because of the smaller training size. But the strong performance of ZeroShot and SPUR demonstrate that LLM-based methods can effectively identify accurate satisfaction/dissatisfaction conversational patterns from limited data.
@@
Importance of Rubric Summarization. Table 2 demonstrates that learning the rubric on each dataset is important for improving the performance on USE. In this experiment, we first use the rubric learned from Bing Copilot (Appendix A.3) in the prompt for MWOZ, SGD and Redial, and evaluate USE performance. Then, we apply the rubrics learned from the target datasets and reevaluate USE performance to gauge how much the Bing Copilot rubrics fail to generalize across tasks. The weighted F1 score in column one shows that rubrics learned on domain-specific data produce an average gain of $13\%$. The numbers of the last two columns are defined by $\widetilde{\mathcal{S}}_{(\cdot)}\setminus\widetilde{\mathcal{S}}_{\text{Bing}}$ and $\widetilde{\mathcal{D}}_{(\cdot)}\setminus\widetilde{\mathcal{D}}_{\text{Bing}}$, which are the set difference between the rubric items in the target and source sets. Values $\geq 0$ indicate that the Rubric Summarization process learns a different set of SAT/DSAT rubrics compared to that of Bing Copilot. This demonstrates that the handcrafted features used by several previous studies (Walker et al., 1997; Bodigutla et al., 2019, 2020) cannot be generalized to different types of conversational systems. However, manually designing rubrics (features) for different conversational systems is ineffective. With our LLM Rubric Summarization process, a targeted set of rubric items can be learned for each task/domain thereby improving USE accuracy.
@@
Figure 3: The average scores for each rubric item w.r.t. click feedback (Like or Dislike). The ‘*’ beside each keyword indicates that the rubric item is significantly correlated with click feedback.
@@
Rubric vs. Click Feedback. Figure 3 shows the correlation between each rubric item and click feedback from users. As discussed in Section 3.3, we ask GPT-4 to generate a label (Yes or No) and a score (0 to 10) for each rubric item in the prompt. The “Yes” label for a rubric item means that the conversational pattern exists in the given conversation, and the score indicates how likely this conversation pattern impacts the overall user satisfaction. The title of each sub-figure in Figure 3 provides a short keyword to summarize the rubric item, and the full descriptions of these keywords are listed in Table 4. The x-axis shows click feedback from users (Like or Dislike). The y-axis shows the average score for each rubric item with respect to the conversations with particular user satisfaction labels. The satisfaction rubric items, which are in the top row, have a higher average score when click feedback is Like. Conversely, the conversations where click feedback is Dislike have higher scores for the dissatisfaction rubric items (bottom row).
@@
From Figure 3, we can see that all twenty rubric items exhibit a significant difference in scores with respect to click feedback. This indicate that the score for each rubric item can be used to improve USE predictions. We conducted a Chi-Square test between the labels of each rubric item and click feedback from users to observe whether these rubric items are useful for USE. The “*” beside each keyword indicates that the rubric item is significantly correlated with the signals provided by clicks.
@@
Figure 4: Satisfaction/Dissatisfaction Pattern Distributions.
Pattern Variance for Different Conversational Systems. Figure 4 reports the satisfaction and dissatisfaction rubric items summarized from the Bing Copilot dataset in the top row, and the bottom row shows the rubric items learned from the MWOZ dataset. Different types of conversational patterns can be observed for the two different conversational systems. Each bar indicates the distribution of the number of times that each rubric item appears in a conversation. Because Bing Copilot is a general-purpose conversational system, the summarized rubric items are general conversational patterns. The detailed description of each Bing Copilot rubric item is shown in Table 4 in Appendix F. In contrast, since MWOZ is a booking chatbot, some satisfaction patterns, e.g. booking confirmation or dissatisfaction patterns and plan adaption, are specific to the booking chatbot. The descriptions for each rubric item learned from the MWOZ dataset are listed in Table 5 in Appendix F.
Similarly, different conversational systems have different service targets, and, therefore, the reasons causing user satisfaction/dissatisfaction are related to the target of the system. Because Bing Copilot is a general-purpose Question-Answering system, the inaccuracy of the information contributes to a large portion of the dissatisfaction. While MWOZ is a booking-reservation system, most of the dissatisfaction comes from lack of proactivity or compromise preference, which means that users have to actively search or choose an option which is less expected.
@@
Figure 5: ROC on Knowlede Distillation from GPT-4.
Knowledge Distillation. Although SPUR can be effectively applied to predict user satisfaction as shown above, since SPUR requires GPT-4 prompting, it is still inefficient to apply USE at web scale (e.g., there have been more than 5 billion conversations in Bing Copilot to date Mehdi . To address this, we propose a knowledge distillation process for each of the rubric items to reduce the cost of the evaluation process. Given the rubric item, we prompt GPT-4 to label another 100K Bing Copilot conversations for knowledge distillation (the label represents whether or not the conversational pattern described by the rubric item appears in the conversation). 80K of the conversations is for training and 20K conversation is for testing. We first calculate an embedding for each conversation (e.g., using OpenAI ada-002) and train a classifier (Logistic Regression) to distill knowledge from GPT-4 (i.e., learn a mapping from embedding to rubric label).
@@
We use the above process to distill knowledge from GPT-4 for one of the satisfaction rubric items (Gratitude) and one of the dissatisfaction rubric items (Negative Feedback). Specifically, we train a Gratitude classifier and a Negative-Feedback classifier. The effectiveness of knowledge distillation is shown in Figure 4(a) and Figure 4(b). A higher AUC metric indicates that the classifier can successfully distill the knowledge from GPT-4 for the given rubric item. We compare the performance of the distilled model with two different embeddings: OpenAI’s text-embedding-ada-002 Ada and multilingual E5 E (5). As a baseline we compare to an embedding-based sentiment classifier: XLM-roBERTa XLM-roBERTa .
Figure 6: Distributions of Feedbacks.
@@
Feedback Distributions. After learning the two textual feedback classifiers, we deploy them to a production environment and seek to understand whether they provide different coverage compared to explicit click feedback (i.e.,“Like” or “Dislike”). Figure 6 reports the distribution of the two types of feedback from one week in production. “Textual” feedback records the proportion of conversations that have true labels predicted by the Gratitude classifier (Textual Like) or by the Negative-Feedback classifier (Textual Dislike). Instead of reporting absolute numbers, we report results relative to the proportion of click feedback we observe in the data. Figure 6 shows the relative frequency of click vs. textual feedback. We can observe that users give more positive feedback through clicks and more negative feedback through their utterances. This also demonstrates the importance of mining conversational SAT/DSAT patterns via SPUR.
@@
Rubric as Features. Finally, we seek to understand if combining the rubrics with conversation text embeddings can produce better results using the model proposed in Appendix E. We use Bing Copilot dataset with 100K conversations for this experiment. This experiment varies the training size from $400$ to $90K$ of the data and $10K$ of the data is for testing. The results in Figure 7 indicate that SPUR provides the best F1 results for smaller training set sizes. As the training set size increases, the F1 scores of the proposed Lin-ada (linear regression with OpenAI ada-002 embeddings) and SPUR-Lin-ada (SPUR rubrics and linear regression with OpenAI ada-002 embeddings) both improve compared to our SPUR model and the SOTA embedding ASAP baseline. The results demonstrate that adding the SPUR metrics to the feature vector consistently provides additional USE signals that are not captured by the conversation embeddings. Note, due to the prohibitive cost, we did not retrain SPUR for larger training set sized above 10,000 samples. Thus, the orange dashed line from 10K to 90K training samples indicates the SPUR F1 score for the test set if we only trained with 10K samples.
@@ 5 Conclusion and Limitations
Limitations. Although SPUR outperforms baseline models with limited training sets, an important factor, the framework is costly if the goal is to estimate user satisfaction at the scale of millions of conversations. Although we have proposed a method to distill knowledge from GPT-4, a thorough study is needed to show the robustness of this approach. In future work, we will focus on SPUR efficiency to reduce its cost at scale.
@@ A.3 User Satisfaction Estimation Prompt
Appendix B Baseline Methods
The following models are used as baselines for comparison of the performance of the proposed SPUR model:
  1. 1.
    Lin-ada: Linear regression model with ada-002 embedding [^2] https://platform.openai.com/docs/guides/embeddings/what-are-embeddings
  2. 2.
  3. 3.
  4. 4.
    Zero shot: prompt GPT-4 directly to score conversations for user satisfaction.
  5. 5.
    Few shot: prompt GPT-4 directly to score conversations, include 2 examples of labeled conversations to guide GPT-4 to determine user satisfaction.
  6. 6.
Appendix C Labeling Adjustment for the Open Data
@@
Appendix D Experiment Setup
We prompt whole process of SPUR with GPT-4 and SPUR-Lin-ada is trained on a NVIDIA A100 instance. Every experiment run one time but with large testing size (80% is testing). The hyperparameters are listed as follows:
  • The number of top-k SAT or DSAT pattern for a conversation is $3$
  • The batch size for minibatch is $100$ SAT/DSAT patterns
  • The number of items for satisfaction rubric and dissatisfaction rubric is $10$
Appendix E User Satisfaction Model
@@
The proposed model is depicted in Figure 8. On the left, the conversations are projected into an embedding space using the GPT-3 Ada-002 embeddings. In parallel, the 20 LLM rubrics are computed using the GPT-4-32K LLM on the right. The 1536 dimension conversation embedding vector is concatenated with the 20 SPUR rubric scores to form the final feature vector which is then input to a model such as Linear Regression, Logistic Regression, or a DNN. The output of the model is the final predicted USE score.
@@
Appendix F Usage of AI Assistants
@@
Name
Description
Name
Description
Gratitude
The user thanks or compliments the AI agent for its help, quality, performance, or abilities.
Repetition
The user repeats their query or request multiple times.
Positive Feedback
The user expresses positive emotions or evaluations using words, phrases, punctuation marks, or emoticons.
Errors
The user points out an error, inconsistency, or inaccuracy in the AI’s output or information and does not receive any acknowledgment or apology from the agent.
Engagement
The user engages in a diverse and lengthy conversation with the AI agent, covering multiple topics or domains.
Negative Feedback
The user uses a negative tone or words to express frustration, disappointment, anger, or disrespect towards the AI agent.
Follow-up
The user asks follow-up questions or requests more information from the AI agent that show curiosity and interest in learning more.
Topic Switch
The user changes their topic or query abruptly.
No Frustration
The user does not express any negative emotion toward the AI agent’s responses throughout the conversation.
Lack Visualization
The user does not receive any visual output from the AI agent when they expect images, links, charts, etc.
Suggestion Acceptance
The user accepts or follows the AI agent’s suggestions, recommendations, and feedback without hesitation, resistance, or challenging it.
No Engagement
The user does not engage with the AI agent’s questions, comments, suggestions, feedback requests, etc.
Personal Details
The user initiates or continues a personal conversation with the AI agent by sharing details about themselves or asking how it is doing.
Irrelevant Information
The user receives a generic, vague, irrelevant answer from the AI agent that does not address their specific needs, goals, or preferences.
Task Request
The user requests specific tasks from the AI agent that match its domain and scope of knowledge, abilities, skills, and expertise.
Complex Answer
The user receives a long and complex answer from the AI agent that may be overwhelming, confusing, or too technical for them.
Correction
The user corrects some of the AI agent’s mistakes, guesses, errors, or misunderstandings in a cooperative, trusting, respectful, and polite manner.
Sudden End
The conversation ends abruptly without fulfilling, completing, or addressing the initial request, problem, task, or goal.
Learning
The user enjoys, appreciates, and learns from different formats, styles, modes, and media of outputs and services, as well as information provided, explained, and generated by the AI agent.
Lack Diversity
The user expects a more interactive, engaging, personalized, humorous, and creative response from the AI, rather than a generic, pre-written, factual, technical, verbose one.
@@
Name
Description
Name
Description
Gratitude
The user thanks the AI agent for its service, indicating gratitude and appreciation.
Repetition
The user repeats their query or request multiple times.
Booking Acceptance
The user accepts the AI agent’s suggestions or bookings without asking for changes or alternatives, implying trust and satisfaction.
Lack Feedback
The user does not receive any confirmation or feedback from the AI after making requests, asking questions, or providing information, leading to uncertainty and confusion.
No Frustration
The user does not express any frustration, confusion, or dissatisfaction with the AI agent’s responses or queries throughout the conversation.
Irrelevant Information
The user receives irrelevant or incomplete information from the AI that does not align with their queries or expectations, which shows a lack of understanding or flexibility.
Follow-up
The user asks questions about the information or options provided by the AI agent, showing interest and engagement.
Ignored or Misunderstood
The user feels ignored or misunderstood by the AI as it does not answer some of their questions, acknowledge their inputs, or provide any clarification.
Booking Confirmation
The user confirms their booking details or information with a positive expression, showing agreement and happiness.
Compromise Preference
The user has to compromise on their desired options or criteria because of limited availability or mismatched recommendations from the AI.
Farewell
The user ends the conversation with a polite farewell and no complaints or requests for further assistance.
Lack of Proactivity
The user has to ask basic questions about features or details that the AI should have provided upfront.
Cooperation
The user follows the AI agent’s guidance and prompts without hesitation or objection, indicating acceptance and cooperation.
Plan Adaption
The user changes their mind about something they previously requested or agreed upon (e.g., location preference) without giving a clear reason.
Clarification
The user specifies their preferences or constraints clearly and specifically, showing confidence and comfort in communicating with the AI agent.
Topic Switch
The user switches to a different topic without closing the previous one.
Request Fulfillment
The user receives relevant and helpful information from the AI agent that matches their requests, such as phone number, price, etc.
Failed bookings
The user experienced several failed bookings and received inconsistent information from the AI about availability.
Request Flexibility
The user is able to change their query or ask for different types of information without encountering any errors or misunderstandings from the AI agent.
False Confirmation
The user was misled by the AI’s confirmation messages, which turned out to be false.

### v2

@@ 4.2 Dataset
P R F1 P R F1 P R F1 P R F1
Lin-ada 74.0 72.5 73.3 48.0 24.0 29.1 53.9 34.9 39.4 56.0 27.7 33.6
USDA 49.7 53.3 47.3 38.1 50.7 35.3 66.1 66.3 61.3 56.1 56.8 48.3
ASAP 66.0 70.1 58.4 51.2 56.1 52.5 64.8 69.8 66.3 60.0 63.6 58.3
Reward 43.6 52.0 52.7 63.0 47.4 40.7 65.3 66.9 58.6 44.1 57.7 48.2
ZeroShot 75.5 73.7 68.3 66.2 52.2 53.6 75.3 70.8 71.9 71.2 58.2 57.0
FewShot 63.8 68.8 61.9 68.4 47.4 44.8 67.3 69.9 66.2 41.0 62.2 49.4
RQ 57.7 69.6 57.7 33.3 52.5 38.1 49.6 67.3 54.4 40.6 63.6 49.5
SPUR 76.3 77.2 75.4 65.7 61.6 59.0 73.7 74.1 72.6 68.4 68.7 66.3
4.3 USE under Few-Shot Setting.
Table 1 shows the performance of each model trained with a small number of training examples. The performance scores are the average of five runs in different train/test splits. The performance metrics are weighted based on the label distributions due to the data imbalance in the different datasets. The training set sizes are shown beside the name of each dataset, and the remaining $80\%$ of the data is used for testing. The number of items in the satisfaction and dissatisfaction rubrics is ten, respectively. Three task-oriented datasets have larger training sizes because we want to ensure that there are at least ten conversations with satisfaction labels and ten conversations with dissatisfaction labels to derive SPUR’s rubrics.
The performance difference between ZeroShot and SPUR lies in that the learned rubrics can provide better guidance for LLMs to determine user satisfaction. Comparing the performance between RQ and SPUR in Table 1, the effectiveness of the rubrics can be observed. Prompting with learned rubrics can provide guidance specific to a dataset for LLMs than prompting with a set of manually selected features Bodigutla et al. (2020) used by all datasets. On the other hand, FewShot has worse performance compared to other methods because the examples provided in the prompt cannot cover many types of satisfaction/dissatisfaction conversational patterns, and the decision is usually biased by the examples provided in the prompt.
The performance of the Reward model (reward deberta, ) validates our hypothesis that Reward models used for RLHF is not a good proxy for scoring user satisfaction. Because Reward models are usually trained with auxiliary human feedback, this reward is not learned from the perspective of the user who was involved in the conversation with the AI agent (Kirk et al., 2023).
Embedding methods perform worse than SPUR in Table 1. Due to the smaller training size, embedding methods cannot generalize well, particularly when the class is imbalanced. They usually have lower weighted F1 scores because they cannot accurately classify the conversations from the less likely classes. However, the strong performance of ZeroShot and SPUR demonstrate that LLM-based methods can effectively identify accurate satisfaction/dissatisfaction conversational patterns from limited data.
@@
Table 2 demonstrates that learning the rubric on each dataset is important for improving the performance on USE. In this experiment, we first use the rubric learned from Bing Copilot (Appendix A.3) in the prompt for MWOZ, SGD and Redial, and evaluate USE performance. Then, we apply the specialized rubrics learned from the target datasets and reevaluate USE performance to gauge how much the Bing Copilot rubrics fail to generalize across tasks. The weighted F1 scores in the first column show that rubrics learned on domain-specific data produce an average gain of $13\%$. The last two columns show the set difference between the rubric items in the target and source sets, i.e., $\widetilde{\mathcal{S}}_{(\cdot)}\setminus\widetilde{\mathcal{S}}_{\text{Copilot}}$ and $\widetilde{\mathcal{D}}_{(\cdot)}\setminus\widetilde{\mathcal{D}}_{\text{Copilot}}$. Values $\geq 0$ indicate that the Rubric Summarization process learns a different set of SAT/DSAT rubrics compared to that of Bing Copilot. This demonstrates that the handcrafted features used by several previous studies (Walker et al., 1997; Bodigutla et al., 2019, 2020) are unlikely to generalize across different types of conversational systems. At the same time, manually designing rubrics (features) for each different conversational systems is time consuming and likely to be ineffective. With our LLM Rubric Summarization process, a targeted set of rubric items can be learned for each task/domain, thereby improving USE accuracy.
@@
Figure 3: The average scores for each rubric item w.r.t. thumb feedback (Like or Dislike). The ‘*’ beside each keyword indicates that the rubric item is significantly correlated with thumb feedback.
4.5 Rubric vs. Thumb Feedback.
@@
Figure 3 shows the correlation between each rubric item and thumb feedback from users. As discussed in Section 3.3, we ask GPT-4 to generate a label (Yes or No) and a score (0 to 10) for each rubric item in the prompt. The “Yes” label for a rubric item means that the conversational pattern exists in the given conversation, and the score indicates how likely this conversation pattern impacts the overall user satisfaction. The title of each sub-figure in Figure 3 provides a short keyword to summarize the rubric item, and the full descriptions of these keywords are listed in Table 4. The x-axis shows thumb feedback from users (Like or Dislike). The y-axis shows the average score for each rubric item with respect to the conversations with particular user satisfaction labels. The satisfaction rubric items, which are in the top row, have a higher average score when thumb feedback is Like. Conversely, the conversations where thumb feedback is Dislike have higher scores for the dissatisfaction rubric items (bottom row).
@@
From Figure 3, we can see that all twenty rubric items exhibit a significant difference in scores with respect to thumb feedback. This indicates that the score for each rubric item can be used to improve USE predictions. We conducted a Chi-Square test between the labels of each rubric item and thumb feedback from users to observe whether these rubric items are useful for USE. The “*” beside each keyword indicates that the rubric item is significantly correlated ($p<0.05$?) with the signals provided by thumb feedback.
4.6 Pattern Variance for Different Conversational Systems.
Figure 4 reports the satisfaction and dissatisfaction rubric items summarized from the Bing Copilot dataset in the top row, and the bottom row shows the rubric items learned from the MWOZ dataset. Different types of conversational patterns can be observed for the two different conversational systems. Each bar indicates the distribution of the number of times that each rubric item appears in a conversation. Because Bing Copilot is a general-purpose conversational system, the summarized rubric items are general conversational patterns. The detailed description of each Bing Copilot rubric item is shown in Table 4 in Appendix E. In contrast, since MWOZ is a booking chatbot, some satisfaction patterns, e.g. booking confirmation or dissatisfaction patterns and plan adaption, are specific to the booking chatbot. The descriptions for each rubric item learned from the MWOZ dataset are listed in Table 5 in Appendix E.
@@
Figure 4: Satisfaction/Dissatisfaction Conversational Pattern Distributions.
Similarly, different conversational systems have different service targets, and therefore, the reasons causing user satisfaction or dissatisfaction are related to the target of the system. Because Bing Copilot is a general-purpose question-answering system, inaccurate information contributes to a larger portion of dissatisfaction. While MWOZ is a booking-reservation system, more of the dissatisfaction is due to a lack of proactivity or a compromise in preference, which means that users have to actively search or choose an option that is less preferred.
4.7 Knowledge Distillation.
Although SPUR can be effectively applied to predict user satisfaction as shown above, since SPUR requires GPT-4 prompting, it is still inefficient to apply USE at web scale (e.g., there have been more than 5 billion conversations in Bing Copilot to date Mehdi . To address this, we propose a knowledge distillation process for each of the rubric items to reduce the cost of the evaluation process. Given the rubric item, we prompt GPT-4 to label a set of conversations for training (the label represents whether or not the conversational pattern described by the rubric item appears in the conversation). Then we calculate an embedding for the conversation (e.g., using OpenAI ada-002) and train a classifier (Logistic Regression) to distill GPT-4 knowledge (i.e., learn a mapping from embedding to rubric label).
@@
Figure 5: ROC on Knowledge Distillation from GPT-4.
@@
We use the above process to distill knowledge from GPT-4 for one of the satisfaction rubric items (Gratitude) and one of the dissatisfaction rubric items (Negative Feedback). Specifically, we train a Gratitude classifier and a Negative-Feedback classifier. The effectiveness of knowledge distillation is shown in Figure 5(a) and Figure 5(b). A higher AUC metric indicates that the classifier can successfully distill the knowledge from GPT-4 for the given rubric item. We compare the performance of the distilled model with two different embeddings: OpenAI’s text-embedding-ada-002 Ada and multilingual E5 E (5). As a baseline we compare to an embedding-based sentiment classifier: XLM-roBERTa XLM-roBERTa . The results show that ada-002 is the most effective text embedding model for knowledge distillation, so we use that in the experiments below.
@@
Feedback Distributions. After learning the two textual feedback classifiers, we deploy them to a production environment and seek to understand whether they provide different coverage compared to explicit thumb feedback (i.e.,“Like” or “Dislike”). Figure 6 reports the distribution of the two types of feedback from one week in production. “Textual” feedback records the proportion of conversations that have true labels predicted by the Gratitude classifier (Textual Like) or by the Negative-Feedback classifier (Textual Dislike). Instead of reporting absolute numbers, we report results relative to the proportion of thumb feedback we observe in the data. Figure 6 shows the relative frequency of thumb vs. textual feedback. We can observe that users give more positive feedback through thumb feedback and more negative feedback through their utterances. This also demonstrates the importance of mining conversational SAT/DSAT patterns via SPUR.
Figure 6: Distributions of Click and Textual Feedback.
4.8 Rubrics as Features
@@
Finally, we seek to understand if combining the rubrics with conversation text embeddings can produce better results using the model proposed in Appendix D. We use Bing Copilot dataset with 100K conversations for this experiment. This experiment varies the training size from $400$ to $90K$ of the data and $10K$ of the data is for testing. The results in Figure 7 indicate that SPUR provides the best F1 results for smaller training set sizes. As the training set size increases, the weighted F1 scores of the SPUR-Lin-ada (SPUR rubrics and linear regression with OpenAI ada-002 embeddings) improves compared to our SPUR model and the SOTA embedding ASAP baseline. The results demonstrate that adding the SPUR metrics to the feature vector consistently provides additional USE signals that are not captured by the conversation embeddings. Note, due to the prohibitive cost, we did not retrain SPUR for larger training set sized above 10,000 samples. Thus, the orange dashed line from 10K to 90K training samples indicates the SPUR F1 score for the test set if we only trained with 10K samples.
@@ 5 Conclusion and Limitations
Limitations. Although SPUR outperforms baseline models with limited training sets, an important factor, the framework is costly if the goal is to estimate user satisfaction at the scale of millions of conversations. We have proposed a method to distill knowledge from GPT-4, but a thorough study is needed to show the robustness of this approach. In future work, we will focus on the scalability issues of SPUR to reduce its cost at a larger scale.
@@ A.3 User Satisfaction Estimation Prompt
Appendix B Labeling Adjustment for the Open Data
@@
Appendix C Experiment Setup
We use GPT-4 for the entire process of training and evaluating SPUR, and SPUR-Lin-ada is trained and tested on an NVIDIA A100 instance. Every experiment runs one time but with a large testing size (80% is used for testing). The hyperparameters are listed as follows:
  • The number of top-k SAT or DSAT patterns for a conversation is $3$.
  • The batch size for each minibatch is $100$ SAT/DSAT patterns.
  • The number of items for the satisfaction rubric and dissatisfaction rubric is $10$.
Appendix D User Satisfaction Model
@@
The proposed model is depicted in Figure 8. On the left, the conversations are projected into an embedding space using the GPT-3 Ada-002 embeddings. In parallel, the 20 LLM rubric itmes are computed using the GPT-4-32K LLM on the right. The 1536-dimension conversation embedding vector is concatenated with the 20 SPUR rubric scores to form the final feature vector which is then input to a model such as Linear Regression, Logistic Regression, or a DNN. The output of the model is the final predicted USE score.
@@
Appendix E Usage of AI Assistants
@@
Name Description Name Description
Gratitude The user thanks or compliments the AI agent for its help, quality, performance, or abilities. Repetition The user repeats their query or request multiple times.
Positive Feedback The user expresses positive emotions or evaluations using words, phrases, punctuation marks, or emoticons. Errors The user points out an error, inconsistency, or inaccuracy in the AI’s output or information and does not receive any acknowledgment or apology from the agent.
Engagement The user engages in a diverse and lengthy conversation with the AI agent, covering multiple topics or domains. Negative Feedback The user uses a negative tone or words to express frustration, disappointment, anger, or disrespect towards the AI agent.
Follow-up The user asks follow-up questions or requests more information from the AI agent that show curiosity and interest in learning more. Topic Switch The user changes their topic or query abruptly.
No Frustration The user does not express any negative emotion toward the AI agent’s responses throughout the conversation. Lack Visualization The user does not receive any visual output from the AI agent when they expect images, links, charts, etc.
Suggestion Acceptance The user accepts or follows the AI agent’s suggestions, recommendations, and feedback without hesitation, resistance, or challenging it. No Engagement The user does not engage with the AI agent’s questions, comments, suggestions, feedback requests, etc.
Personal Details The user initiates or continues a personal conversation with the AI agent by sharing details about themselves or asking how it is doing. Irrelevant Information The user receives a generic, vague, irrelevant answer from the AI agent that does not address their specific needs, goals, or preferences.
Task Request The user requests specific tasks from the AI agent that match its domain and scope of knowledge, abilities, skills, and expertise. Complex Answer The user receives a long and complex answer from the AI agent that may be overwhelming, confusing, or too technical for them.
Correction The user corrects some of the AI agent’s mistakes, guesses, errors, or misunderstandings in a cooperative, trusting, respectful, and polite manner. Sudden End The conversation ends abruptly without fulfilling, completing, or addressing the initial request, problem, task, or goal.
Learning The user enjoys, appreciates, and learns from different formats, styles, modes, and media of outputs and services, as well as information provided, explained, and generated by the AI agent. Lack Diversity The user expects a more interactive, engaging, personalized, humorous, and creative response from the AI, rather than a generic, pre-written, factual, technical, verbose one.
@@
Name Description Name Description
Gratitude The user thanks the AI agent for its service, indicating gratitude and appreciation. Repetition The user repeats their query or request multiple times.
Booking Acceptance The user accepts the AI agent’s suggestions or bookings without asking for changes or alternatives, implying trust and satisfaction. Lack Feedback The user does not receive any confirmation or feedback from the AI after making requests, asking questions, or providing information, leading to uncertainty and confusion.
No Frustration The user does not express any frustration, confusion, or dissatisfaction with the AI agent’s responses or queries throughout the conversation. Irrelevant Information The user receives irrelevant or incomplete information from the AI that does not align with their queries or expectations, which shows a lack of understanding or flexibility.
Follow-up The user asks questions about the information or options provided by the AI agent, showing interest and engagement. Ignored or Misunderstood The user feels ignored or misunderstood by the AI as it does not answer some of their questions, acknowledge their inputs, or provide any clarification.
Booking Confirmation The user confirms their booking details or information with a positive expression, showing agreement and happiness. Compromise Preference The user has to compromise on their desired options or criteria because of limited availability or mismatched recommendations from the AI.
Farewell The user ends the conversation with a polite farewell and no complaints or requests for further assistance. Lack of Proactivity The user has to ask basic questions about features or details that the AI should have provided upfront.
Cooperation The user follows the AI agent’s guidance and prompts without hesitation or objection, indicating acceptance and cooperation. Plan Adaption The user changes their mind about something they previously requested or agreed upon (e.g., location preference) without giving a clear reason.
Clarification The user specifies their preferences or constraints clearly and specifically, showing confidence and comfort in communicating with the AI agent. Topic Switch The user switches to a different topic without closing the previous one.
Request Fulfillment The user receives relevant and helpful information from the AI agent that matches their requests, such as phone number, price, etc. Failed bookings The user experienced several failed bookings and received inconsistent information from the AI about availability.
Request Flexibility The user is able to change their query or ask for different types of information without encountering any errors or misunderstandings from the AI agent. False Confirmation The user was misled by the AI’s confirmation messages, which turned out to be false.